OpenStudy (anonymous):

\[10011_{2}\] x \[11_{2}\]

OpenStudy (anonymous):

57

OpenStudy (amistre64):

16 8 4 2 1 2 1 ---------- * --- = 19 * 3 = 30 + 27 = 57 1 0 0 1 1 1 1